OpenStudy (anonymous):

A cylinder labeled X has a height of 18 inches and a radius of 21 inches. Another cylinder labeled Y has height of 90 inches and radius of 105 inches. The volume of cylinder Y is _______ times the volume of cylinder X @Hero

OpenStudy (anonymous):

@dpaInc

hero (hero):

I guess the note taking techniques didn't work out for you

OpenStudy (anonymous):

I know the answer but I am not sure

hero (hero):

Hint: Apply the formula for Volume of a Cylinder

OpenStudy (anonymous):

is it 5 times bigger?

hero (hero):

Did you apply the formula, or are you guessing?

OpenStudy (anonymous):

I applied the formula

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

@Calcmathlete

hero (hero):

I think you were guessing rather than using the formula

OpenStudy (anonymous):

let me redo, then

hero (hero):

2nd Hint: \[\frac{\text{Volume of Y}}{\text{Volume of X}}\]

OpenStudy (anonymous):

v=pi*r^2*h v=3.14*21^2*18 v=24925.32

hero (hero):

\[\frac{\pi (105)^2 90}{\pi (21)^2 18}\]

hero (hero):

Simplify that

OpenStudy (anonymous):

v=pi*r^2*h v=3.14*105^2*90 v=3115665

OpenStudy (anonymous):

=125 times

hero (hero):

There we go

OpenStudy (anonymous):

so its really 125 times bigger?

hero (hero):

I know this comes as a surprise to you, but yes

OpenStudy (anonymous):

ok got it, I could have used the ratio formula as well! thanx so much!

OpenStudy (anonymous):

thanx so much @Calcmathlete

hero (hero):

I thought I was the one who helped you

OpenStudy (lukecrayonz):

God I love you @Hero